# Repurposing 2D Diffusion Models with Gaussian Atlas for 3D Generation

### v1

@@ 2 Related Work
3D generation with Gaussian splatting. Unlike 2D images, 3D objects are more difficult to generate due to the additional dimension and geometric constraints [28, 15]. Among all 3D generation methods [47, 34, 16, 4, 48, 51, 18], there is a line of work that utilizes diffusion models to generate 3D Gaussians, which is closely related to this paper. GSD [29] introduced rendering guidance to constrain the sampling of 3D Gaussians with 2D observations. L3DG [35] is another example, which embeds 3D Gaussians onto a dense latent grid and learns a diffusion model in the latent space for generation. Moreover, GVGen [10] and GaussianCube [55] achieve 3D Gaussian diffusion by transforming sparsely located Gaussians into more structured 3D volumes. GVGen fits 3D Gaussians directly on a volume with offsets, while GaussianCube applies Optimal Transport to move unorganized 3D Gaussians to vertices of a predefined 3D grid. Most recently, Gaussian Anything [17] adopts a two-step approach that generates Gaussians by auto-encoding latents in the point cloud space. In this work, we propose a novel structuralization approach that enables 2D transformation for 3D Gaussians, to leverage the power of well pretrained 2D diffusion models and unify the architecture of both 2D and 3D generation.
@@
2D representations of 3D content. Representing 3D objects as 2D planes is not new. A major line of research encodes 3D structures using implicit Triplanes [3]. NFD [41] was one of the first attempts to train 2D diffusion models to generate Triplanes, from which neural fields can be derived to reconstruct 3D objects. To enhance 3D coherence, CRM [48] integrates multi-view observations into the Triplane generation process. InstantMesh [51] and Instant3D [18] follow a similar design but employ more sophisticated models to improve Triplane synthesis. Instead of generating neural fields, TriplaneGaussian [58] decodes 3D Gaussian attributes directly from generated Triplanes. More recently, DiffGS [57] has refined the TriplaneGaussian paradigm by learning a mapping between pre-fitted 3D Gaussians and Triplanes via a latent diffusion model. Several works share motivations similar to ours. For example, PI3D [20] fine-tunes a pretrained text-to-image generation model to produce Triplanes represented as “pseudo-images,” while HexaGen3D [27] leverages pretrained image diffusion models for improved Triplane generation. However, their fine-tuned diffusion models exhibit limited capacity, and an additional minutes-long refinement process is typically required to enhance the 3D generation.
@@
In addition to Triplane-based representations, Splatter Image [44] trains a 2D network to infer pixel-wise Gaussian attributes from single images; Omages [53] trains 2D diffusion models to fit 2D UV maps of the geometry and materials of 3D objects, ultimately reconstructing a textured 3D mesh. However, Omegas supports only class-conditioned generations of a limited pre-defined categories; GIMDiffusion [5] adopts a similar approach by representing the surface of 3D objects with geometry images [8] and leveraging 2D diffusion models in the UV space. However, their generated UV maps are sparse, leading to artifacts at the seams of the assembled 3D mesh. In contrast, our method introduces a dense 2D representation of 3D Gaussians, thereby mitigating these limitations.
@@ 3 GaussianVerse
With the above fitting strategy, we optimize per-object 3D Gaussians by minimizing photometric losses against multi-view RGB renderings:
@@ 4 Formulating 3D Gaussians as 2D Atlas
However, sparse and unstructured Gaussians in 3D space cannot be directly passed to 2D models, which require inputs $\mathbf{X}$ to have: (i) only 2 spatial dimensions $W$ and $H$ such that $\mathbf{X}\in\mathbb{R}^{W\times H\times\cdot}$; (ii) valid “pixels” at each vertex of a dense 2D grid; and (iii) values within a specific distribution, either $\mathbf{X}\in[-1,1]$ for the VAE or $\mathbf{X}\sim\mathcal{N}(0,1)$ for the latent UNet. To make 3D Gaussians compatible with 2D diffusion models, we propose Gaussian Atlas, a 2D representation of 3D Gaussians.
@@
Specifically, our goal is to find a mapping function $\mathcal{M}(\cdot)$ that transforms the 3D positions $\{\mathbf{x}\in\mathbb{R}^{3}\}$ of 3D Gaussians to 2D planar coordinates $\{\mathbf{\hat{x}}\in\mathbb{R}^{2}\}$: ($\mathcal{M}(\{\mathbf{x}\})\rightarrow\{\mathbf{\hat{x}}\}$. A natural approach is to parameterize such a function $\mathcal{M}$ as neural networks which can be optimized toward 2D and 3D consistency [56]. However, this approach requires repeated training of $\mathcal{M}$ on different objects which is not only time consuming but also makes the mapping process inconsistent between objects given that the heuristics for 2D flattening may be different for different geometries. As a result, diffusion models are not able to capture the irregular patterns and fail to generate meaningful contents. Experiments are provided in section 6.2 to support this claim. To this end, we seek a more coherent transformation between 3D and 2D with a simple and deterministic mapping function.
@@
After positioning the 3D Gaussians on the surface of a unit sphere, we employ the classic equirectangular projection [49] as our function $\mathcal{M}$ to obtain flattened 2D coordinates $\{p_{i}\in\mathbb{R}^{2}\}$ of the 3D Gaussians. Lastly, in the plane offsetting stage, we apply another OT to map $\{p_{i}\}$ to the vertices $\{q_{i}\in\mathbb{R}^{2}\}$ of a square 2D grid with spatial size $\sqrt{N}\times\sqrt{N}$, further reducing sparsity. In particular, since now $\mathcal{M}$ is a deterministic function, the mapping between $\{p_{i}\}$ and $\{q_{i}\}$ remains identical between different 3D objects. This consistency allows us to perform OT in the last stage only once and reuse the computed indices for all objects. We refer to the final grid-like 2D representation of 3D Gaussians as Gaussian Atlas, with each atlas $\mathbf{X}$ in the shape of $\sqrt{N}\times\sqrt{N}\times(||\mathbf{x}-\mathbf{s}||+||\mathbf{c}||+||\mathbf{o}||+||\mathbf{s}||+||\mathbf{r}||)$ entailing all attributes of 3D Gaussians.
@@
Figure 5: Qualitative Comparisons. Our 3D generations exhibit the highest quality, minimal artifacts, and the best alignment with text prompts. In contrast, DreamGaussian [45], LGM [46], and TriplaneGaussian [58] struggle to produce natural and complete 3D assets, while GaussianCube [55] fails to capture key phrases in text prompts, such as “on a plate” and “pink rooftop.”
@@
where $\mathcal{L}_{rgb},\mathcal{L}_{mask}$ are L1 loss on RGB renderings and accumulated opacity maps and $\mathcal{L}_{lpips}$ is the perceptual loss. $\lambda$s are the corresponding loss weights. The overall pipeline is outlined in Figure 4.
@@ 6 Experiments
In this section, we benchmark and evaluate the effectiveness of our proposed methods on zero-shot text-to-3D generation, one of the most fundamental tasks in generative 3D modeling.
@@
We initialize the 2D UNet $\mathcal{F}$ from a LD checkpoint pretrained on LAION [39] and use CLIP [33] to encode text prompts. The UNet is then fine-tuned with classifier-free guidance by randomly zeroing out the text encodings with a probability of 20%. We employ the AdamW optimizer [21] with a learning rate of $5\times 10^{-5}$ and a batch size of 64. Exponential Moving Average (EMA) and mixed precision training are enabled for stable and efficient training. For fair comparisons with the baseline method [55], only approximately 130K atlases from GaussianVerse were used for training. The paired text prompts were from [25] and the multi-view references were from [59]. The LD UNet was finetuned for 1M steps on 8 A100 GPUs.
For inference, we start from 2D random noise in the same shape as a Gaussian atlas and generate clean samples via reverse diffusion using the adaptive DPMsolver+++++ + [22] with a guidance scale of 3.5. A 3DGS sample can be generated and rendered in less than 5 seconds.
@@ 6.1 Results
Qualitative results. We present zero-shot 3D generation results in Figure 5. Compared to DreamGaussian and LGM, our generations exhibit high visual fidelity without excessive details. In contrast to TriplaneGaussian and GaussianCube, our method aligns closely with the text prompts, reflecting more precise prompt conditioning in the generations. According to the additional qualitative results shown in Figure 6, our fine-tuned 2D diffusion model is capable of generating high-quality Gaussian atlases that effectively generalize to 3D information. Consequently, our generated 3D contents are coherent and self-contained, with no seams or artifacts that would typically be observed from UV unwrapping methods. We provide more qualitative results in the supplementary materials.
User studies [^1] The user study was conducted entirely by researchers at Stanford.. The user study was conducted with a diverse group of participants who were asked to assess the overall generation quality corresponding to given text prompts. They performed pairwise comparisons between generations by our method and those from a competing method. We calculated the win rate from the collected $>2,500$ valid responses. As shown in Figure 7, over 65% of users preferred the 3D content generated by our method when compared to the state-of-the-art GaussianCube. The preference rate was even more pronounced against TriplaneGaussian, with 88% of users favoring our method.
User studies. The user study was conducted with a diverse group of participants who were asked to assess the overall generation quality corresponding to given text prompts. They performed pairwise comparisons between generations by our method and those from a competing method. We calculated the win rate from the collected $>2,500$ valid responses. As shown in Figure 7, over 65% of users preferred the 3D content generated by our method when compared to the state-of-the-art GaussianCube. The preference rate was even more pronounced against TriplaneGaussian, with 88% of users favoring our method.
@@ 6.2 Discussions and Ablations
$\bullet$  Learned prior knowledge in text-to-image diffusion models is universally transferable. Without pre-training on large-scale 2D datasets, the model struggles to develop a comprehensive understanding of natural language and content generation when trained solely on 3D data. As shown in Figure 8, we effectively repurpose the pre-trained 2D diffusion model, resulting in significantly faster convergence and higher-quality 3D generation. The qualitative results show that, compared to training from scratch, our approach produces more coherent and structured 3D assets earlier in training. The quantitative results further support this finding, as our repurposed models consistently achieve higher scores in different training steps.
@@
Figure 9: The optimization-based flattening approach results in inconsistent patterns and poses challenges for fine-tuning. We present two objects (a) and their corresponding Gaussian atlases (3D location $\mathbf{x}$ only) obtained using our proposed flattening approach (b) and the optimization-based approach [56] (c), which generates only noisy Gaussians (d) and atlases (e) after training for the same number of steps.
$\bullet$  A pre-trained 2D diffusion model serves as a better initialization for 3D generation. In Figure 10, we plot the weight differences between our fine-tuned UNet and both a randomly initialized UNet and the pre-trained UNet directly from LD. We observe overall small changes in the weights from the pre-trained UNet. Even at the layer with the greatest change, the difference is still 8×8$\times$8 × smaller than that from a random initialization. This indicates that the pre-trained LD weights provide better starting points for 3D generation, facilitating easier convergence to a local minimum.

### v2

@@ 2 Related Work
3D generation with Gaussian splatting. Unlike 2D images, 3D objects are more difficult to generate due to the additional dimension and geometric constraints [29, 15]. Among all 3D generation methods [48, 35, 16, 4, 49, 53, 18], there is a line of work that utilizes diffusion models to generate 3D Gaussians, which is closely related to this paper. GSD [30] introduced rendering guidance to constrain the sampling of 3D Gaussians with 2D observations. L3DG [36] is another example, which embeds 3D Gaussians onto a dense latent grid and learns a diffusion model in the latent space for generation. Moreover, GVGen [10] and GaussianCube [57] achieve 3D Gaussian diffusion by transforming sparsely located Gaussians into more structured 3D volumes. GVGen fits 3D Gaussians directly on a volume with offsets, while GaussianCube applies Optimal Transport to move 3D Gaussians to vertices of a predefined 3D grid. Gaussian Anything [17] adopts a two-step approach that generates Gaussians by auto-encoding point cloud latents. Recently, TRELLIS [51] designed a structured representation that supports decoding to various 3D formats. In this work, we propose to transform 3D Gaussians onto 2D atlases, to leverage the power of well pretrained 2D diffusion models and unify the architecture of 2D and 3D generation.
@@
2D representations of 3D content. Representing 3D objects as 2D planes is not new. A major line of research encodes 3D structures using implicit Triplanes [3]. NFD [42] was one of the first attempts to train 2D diffusion models to generate Triplanes, from which neural fields can be derived to reconstruct 3D objects. To enhance 3D coherence, CRM [49] integrates multi-view observations into the Triplane generation process. InstantMesh [53] and Instant3D [18] follow a similar design but employ more sophisticated models to improve Triplane synthesis. Instead of generating neural fields, TriplaneGaussian [60] decodes 3D Gaussian attributes directly from generated Triplanes. More recently, DiffGS [59] has refined the TriplaneGaussian paradigm by learning a mapping between pre-fitted 3D Gaussians and Triplanes via a latent diffusion model. Several works share motivations similar to ours. For example, PI3D [21] fine-tunes a pretrained text-to-image generation model to produce Triplanes represented as “pseudo-images”, while HexaGen3D [28] leverages pretrained image diffusion models for improved Triplane generation. However, their fine-tuned diffusion models exhibit limited capacity, and an additional minutes-long refinement process is typically required to enhance the 3D generation.
@@
In addition to Triplane-based representations, Splatter Image [45] trains a 2D network to infer pixel-wise Gaussian attributes from single images; Omages [55] trains 2D diffusion models to fit 2D UV maps of the geometry and materials of 3D objects, ultimately reconstructing a textured 3D mesh. However, Omegas supports only class-conditioned generations of a limited pre-defined categories; GIMDiffusion [5] adopts a similar approach by representing the surface of 3D objects with geometry images [8] and leveraging 2D diffusion models in the UV space; DiffSplat [19] trains VAE to compress multi-view 3D Gaussians into 2D latents and achieve multi-view generation through a pretrained 2D diffusion model. In contrast, our method introduces a more dense 2D representation of 3D Gaussians, yielding a novel approach for repurposing pretrained 2D diffusion models.
@@ 3 GaussianVerse
We optimize per-object 3D Gaussians by minimizing photometric losses against multi-view RGB renderings:
@@ 4 Formulating 3D Gaussians as 2D Atlas
However, unstructured Gaussians in 3D space cannot be directly passed to 2D models, which require inputs $\mathbf{X}$ to have: (i) only 2 spatial dimensions; (ii) valid “pixels” at each vertex of a dense 2D grid; and (iii) values within a specific distribution, either $\mathbf{X}\in[-1,1]$ for the VAE or $\mathbf{X}\sim\mathcal{N}(0,1)$ for the denoiser. To make 3D Gaussians compatible with 2D diffusion models, we propose Gaussian Atlas, a 2D representation of 3D Gaussians.
@@
Specifically, our goal is to find a mapping function $\mathcal{M}(\cdot)$ that transforms the 3D positions $\{\mathbf{x}\in\mathbb{R}^{3}\}$ of 3D Gaussians to 2D planar coordinates $\{\mathbf{\hat{x}}\in\mathbb{R}^{2}\}$: $\mathcal{M}(\{\mathbf{x}\})\rightarrow\{\mathbf{\hat{x}}\}$. A natural approach is to parameterize such a function $\mathcal{M}$ as neural networks which can be optimized toward 2D and 3D consistency [58]. However, this approach requires repeated training of $\mathcal{M}$ on different objects which is not only time consuming but also makes the mapping process inconsistent between objects given that the heuristics for 2D flattening may be different for different geometries. As a result, diffusion models are not able to capture the irregular patterns and fail to generate meaningful contents. Experiments are provided in section 6.2 to support this claim. To this end, we seek a more coherent transformation between 3D and 2D with a simple and deterministic mapping function.
@@
[…]
@@
Figure 5: Qualitative Comparisons. Our 3D generations exhibit the highest quality, minimal artifacts, and the best alignment with text prompts. In contrast, DreamGaussian [46], LGM [47], and TriplaneGaussian [60] struggle to produce natural and complete 3D assets, while GaussianCube [57] fails to capture key phrases in text prompts, such as “on a plate” and “pink rooftop”.
@@
[…]
@@ 6 Experiments
In this section, we benchmark and evaluate our proposed methods on zero-shot text-to-3D generation, one of the most fundamental tasks in generative 3D modeling.
@@
We initialize the 2D UNet $\mathcal{F}$ from a LD pretrain checkpoint [40] and use CLIP [34] to encode text prompts. The UNet is then fine-tuned with classifier-free guidance by randomly zeroing out the text encodings with a probability of 20%. We employ the AdamW optimizer [22] with a learning rate of $5\times 10^{-5}$ and a batch size of 64. Exponential Moving Average (EMA) and mixed precision training are enabled for stable and efficient training. For fair comparisons with the baseline method [57], only approximately 130K atlases from GaussianVerse were used for training. The paired text prompts were from [26] and the multi-view references were from [61]. The LD UNet was finetuned for 1M steps on 8 A100 GPUs.
For inference, we start from 2D random noise in the same shape as a Gaussian atlas and generate clean samples via reverse diffusion using the adaptive DPMsolver++++ [23] with a guidance scale of 3.5. A 3DGS sample can be generated and rendered in less than 5 seconds.
@@ 6.1 Results
Qualitative results. We present text-to-3D generation results in Figure 5. Compared to DreamGaussian and LGM, our generations exhibit high visual fidelity without excessive details. In contrast to TriplaneGaussian and GaussianCube, our method aligns closely with the text prompts, reflecting more precise prompt conditioning. According to the additional results shown in Figure 6, our fine-tuned 2D diffusion model is capable of generating high-quality Gaussian atlases. Consequently, our generated 3D contents are coherent and self-contained, with no seams or artifacts that would typically be observed from UV unwrapping methods.
Quantitative results. In Table 2, we present the CLIP and VQA scores for comparison methods. Our method outperforms all counterparts while generating the minimum number of Gaussians required. Specifically, it uses only $\frac{1}{2}$ the number of Gaussians and the number of training steps on 3D data compared to GaussianCube, achieving a CLIP score higher by 0.9 and a VQA score higher by 17%. This further supports our claim that pretrained 2D diffusion models can be repurposed for 3D content generation.
User studies [^1] The user study was conducted entirely by researchers at Stanford.. The user study was conducted with diverse participants who were asked to assess the overall generation quality to given text prompts. They performed pairwise comparisons between generations by our method and those from a competing method. We calculated the win rate from >2,500>2,500 valid responses. As shown in Figure 7, over 65% of users preferred the 3D content generated by our method when compared to GaussianCube. The preference rate was even more pronounced against TriplaneGaussian, with 88% of users favoring our method.
@@ 6.2 Discussions and Ablations
$\bullet$  Learned knowledge in text-to-image diffusion models is universally transferable. Without pre-training on large-scale 2D data, the model struggles to develop a comprehensive understanding of natural language and content generation when trained solely on 3D data. As shown in Figure 8, we effectively repurpose the pre-trained 2D diffusion model with faster convergence and higher-quality 3D generation. The qualitative results show that, compared to training from scratch, our approach produces more coherent and structured 3D assets earlier in training. The quantitative results further support this finding, as our repurposed models consistently achieve higher scores in different training steps.
@@
Figure 9: The optimization-based flattening approach results in inconsistent patterns that are hard for fine-tuning. We show two objects (a) and their Gaussian atlases (3D location $\mathbf{x}$ only) obtained using our proposed flattening approach (b) and the optimization-based approach [58] (c), which generates only noisy Gaussians (d) and atlases (e) after training for the same number of steps.
$\bullet$  A pre-trained 2D diffusion model serves as a better initialization for 3D generation. In Figure 10, we plot the weight differences between our fine-tuned UNet and both a randomly initialized UNet and the pre-trained UNet directly from LD. We observe overall small changes in the weights from the pre-trained UNet. Even at the layer with the greatest change, the difference is still $8\times$ smaller than that from a random initialization. This indicates that the pre-trained LD weights provide better starting points for 3D generation, facilitating easier convergence to a local minimum.
@@ 7 Conclusion
Acknowledgments.
Tiange Xiang, Scott Delp, Ehsan Adeli, and Li Fei-Fei’s work at Stanford were partially funded by the NIH Grant R01AG089169 and P41EB027060, Panasonic Holdings Corporation, the Gordon and Betty Moore Foundation, the Jaswa Innovator Award, Stanford HAI, Stanford HAI graduate fellowship, and Stanford Wu Tsai Human Performance Alliance. We thank Fei Jiang and Moustafa Meshry for their insightful feedback on this work.